OpenStudy (anonymous):

Solve for x 8(a+x)=b-a+6x

OpenStudy (texaschic101):

8(a + x) = b - a + 6x -- divide both sides by 8 a + x = 1/8b - 1/8a + 3/4x -- subtract a from both sides x = 1/8b - 1/8a - a + 3/4x -- subtract 3/4x 4/4x - 3/4x = 1/8b - 1/8a - 8/8a 1/4x = 1/8b - 9/8a -- divide by 1/4 x = (1/8b - 9/8a) / (1/4) x = (1/8b - 9/8a) * 4 x = 1/2b - 36/8a x = (1/2)b - (9/2)a

OpenStudy (anonymous):

Ty